the average rainfall of a city from monday to saturday is 0.3cm due to heavy rainfall on sunday the average for the whole week rose to 0.5 cm . how much was the rainfall on sunday.
Answer by Cromlix(4381)

Average rainfall for 6 days = 0.3cm
Total rainfall for 6 days = 6 * 0.3 = 1.8 cm
Average rainfall for 7 days = 0.5
Total rainfall for 7 days = 7 * 0.5 = 3.5 cm
3.5 - 1.8 = 1.7 cm
Thus, rainfall for Sunday = 1.7 cm.
Hope this helps.
:-)